# Carry-Over Lottery Allocation: Practical Incentive-Compatible Drafts

### v1

@@ Abstract
The NBA Draft lottery is designed to promote competitive balance by awarding better draft positions to weaker teams, but it creates incentives to deliberately lose, a practice known as tanking. We propose a draft mechanism that is simultaneously practical, incentive-compatible, and advantages weaker teams. The Carry-Over Lottery Allocation (COLA) Draft Mechanism represents a paradigm shift in evaluating team quality, replacing a single season’s standings with playoff outcomes over multiple years. COLA uses a draft lottery where every non-playoff team receives the same number of lottery tickets, removing incentives to lose additional games after elimination. Lottery tickets that do not win a top draft pick carry over to future lotteries, while playoff success or winning a top pick diminishes a team’s accumulated tickets. Over time, COLA rewards teams with poor long-term performance and less prior draft assistance. By retaining the lottery format, COLA preserves transparency and fan engagement.
Real-world implementation challenges are addressed to demonstrate feasibility, including transitioning from the current system, handling traded draft picks, and accommodating draft classes of varying strength. The most significant challenge occurs in years with exceptionally strong draft classes, where teams may prefer missing the playoffs in order to gain lottery access, violating a foundational assumption: that teams prefer playoff success to lottery participation. We provide a solution to this problem, employing a truth-elicitation mechanism to identify such years and expand lottery eligibility to include as many playoff teams as necessary to preserve anti-tanking incentives.
@@ Introduction
In the current NBA draft system, all non-playoff teams enter a lottery for the top four draft positions, with worse records giving higher chances of winning. Teams that do not win a top pick are then slotted by record, starting with the worst team. This structure creates a trade-off: the league wants to reward weaker teams to maintain competitive balance, but higher lottery odds for the worst teams also increase incentives to lose. Despite repeated adjustments to lottery probabilities, tanking persists because of a fundamental impossibility: Munro and Banchio [7] proved that no draft mechanism based solely on end-of-season records can simultaneously reward poor performance and eliminate incentives to tank. The current system is, mathematically, doomed to fail.
@@
We answer in the affirmative with the proposal of the Carry-Over Lottery Allocation (COLA) Draft Mechanism. COLA achieves incentive compatibility by giving all non-playoff teams the same number of lottery tickets each year, eliminating any advantage from additional losses. COLA rewards the weakest teams through carry-over: lottery tickets that do not win a top draft pick are retained for future lotteries, while playoff success or winning a top pick diminishes a team’s accumulated tickets.
COLA is also practical to implement. It uses a familiar draft with lottery, requiring minimal changes to existing league procedures and fan expectations. Traded draft picks need special handling, but COLA solves this by excluding traded picks from lottery eligibility entirely. The mechanism also accommodates variations in draft class strength. For a weak class, a team might view a lottery win as undesirable due to a loss of tickets in future lotteries, so there is an option to skip the lottery for a cost. For a strong draft class, many teams might prefer to be in the lottery instead of the playoffs. This violates a key assumption of anti-tanking mechanisms: winning is the primary goal, with draft picks as a secondary consideration. This problem has no solution in existing mechanisms, because if all teams prefer the lottery, no mechanism can both reward weak teams and prevent tanking. To address this, COLA identifies strong draft classes in advance using a Bayesian Truth Serum–based media survey and correspondingly expands lottery access as far as necessary but only when necessary, maintaining incentive compatibility without altering the core structure of the mechanism.
The remainder of this paper proceeds as follows. In Section 2, we review related work, including other draft mechanisms. Section 3 presents the core mechanism: how lottery tickets accumulate, how they are diminished by playoff success or lottery wins, and why this structure ensures that teams never maximize their expected utility by losing. Section 4 examines critical implementation challenges, including the handling of traded draft picks, transitioning from the current system, and addressing potential violations of our preference assumptions. Section 5 presents simulation results that provide further evidence that under COLA all teams have an equal likelihood of success over time.
@@
  • Munro and Banchio [7] propose that the lottery be based on standings at a dynamically-determined point in the season. It meets the Preferencing and Anti-Tanking criteria under standard assumptions, but fails Practicality for two reasons. First, the calculation of the dynamically-determined date is opaque to a casual fan and requires knowledge of private team preferences. Choosing a predetermined early cutoff date may seem to approximate their approach, but if the date is announced and does not take into account teams’ incentives, then it simply moves tanking to earlier in the year. Second, in most seasons there are teams that believe their likelihood of playoff success is so small that they have an incentive to tank as soon as the season starts, in which case there are no standings on which to base the lottery.
  • The WNBA bases the lottery on the combined record of the two most recent seasons. This improves Preferencing for persistently weak teams but fails Anti-Tanking even more severely than the NBA’s system. A loss in one season could affect the draft position for two seasons, giving even playoff teams incentives to underperform.
  • Gold Plan [3] bases the draft on the number of wins that a team earns after playoff elimination. Under the assumption that every team puts forth maximum effort until they are mathematically eliminated, it satisfies Anti-Tanking. However, in practice it is common for teams to tank before they are eliminated. This mechanism can incentivize tanking earlier in the season so that they are eliminated as soon as possible. That way, the team has more time to rack up post-elimination wins. It also disadvantages truly bad teams that cannot win many games even if they are eliminated early. Thus, it fails both Practicality and Preferencing.
@@ A New Paradigm for Sports Drafts
With this paradigm shift, we can design a draft mechanism that meets all five of the desired criteria. The following section presents COLA, a practical mechanism that implements these principles.
@@
We introduce the Carry-Over Lottery Allocation (COLA) Draft Mechanism, named for two of its key features: draft picks are allocated via lottery, and some lottery tickets carry over between years.
@@
Increment Size. We use an annual increment of $\alpha=1000$, but a different value could be used. To maintain incentive compatibility, it is important that a team not gain any benefit from additional losses, but it is not necessary for $\alpha$ to be constant. For example, the increment could be larger for teams that have missed the playoffs for multiple years in a row.
@@
Diminishment Rates. Our choice (100%, 75%, 50%, 25% for draft picks; similar for playoff rounds) prioritizes simplicity. More aggressive rates emphasize recency; gentler rates give historical performance more weight. More aggressive rates on playoff success also increase the likelihood that the line must be moved to maintain incentive compatibility. More complex diminishment rules could be employed. Beyond percentages, diminishment could use fixed amounts, caps/floors, or escalation rules (e.g., doubled rates for teams making three consecutive playoff appearances).
@@
Lottery Scope. Following current NBA practice, the lottery is for the top four picks, but this extends naturally to more positions if needed. The mechanism works identically with any number of lottery positions, though more positions increase the potential for the worst team to receive a poor pick despite having the best odds.
@@
Non-Lottery Pick Assignment. Picks 5-14 could be assigned by current-season standings (matching the existing NBA system) or by lottery index (consistent with COLA’s multi-year evaluation). Standings-based assignment provides familiarity, while lottery-index assignment rewards historically weak teams throughout the draft order, not just through the lottery. The latter also mitigates the incentive to tank for the purpose of retaining a traded draft pick that has protections, but those sorts of trade protections should be banned anyway.
@@
No Lottery At All. If the draft pick order is determined by ordering all teams by highest lottery index (as suggested in the previous variant), then Playoff Track Record and prior drafts would determine the draft order instead of regular season standings. Equal increments, carry-over, and diminishment for playoff success and top draft picks accomplish their goals with or without a lottery. (Diminishment for top picks should remain, even if not determined by lottery. Otherwise, the same team would pick first each year until making the playoffs.) However, the lottery also generates fan interest, so preventing tanking is not its only purpose.
@@
Opt-Out Penalty. The $2\alpha$ cost for skipping weak draft classes balances individual flexibility against league-wide participation. Lower the value to increase team autonomy. Adjust higher or eliminate the option if more participation is desired.
@@
Survey Detail. Our six-point scale on the analyst survey could be expanded to more options for more granularity at the expense of complexity for survey participants. For example, an alternative design could expand the survey to twelve options by allowing each playoff round to be included in the lottery either with or without diminishment for playoff success at that level.
@@
Increments in Strong Draft Years. When the line moves to include playoff teams, we use a binary rule: if eight or more teams remain excluded, all lottery participants receive $\alpha$; if four or fewer remain excluded, nobody receives $\alpha$. A simpler alternative would always grant $\alpha$ to all lottery participants regardless of where the line falls, prioritizing ease of explanation over maximizing Preferencing for Quality in strong draft years.
@@
Traded Picks Winning the Lottery. Under COLA, traded draft picks are excluded from lottery eligibility. An alternative design would allow traded picks to enter the lottery, with lottery index diminishment applied to both the pick holder and the original team if the pick wins. This would introduce a third class of traded draft picks: protected 1–4, unprotected but lottery-ineligible, and unprotected lottery-eligible picks. While this variant improves trade-market liquidity and gives teams more flexibility, it is substantially more complex and creates the opportunity for systematic mismanagement of long-run lottery risk by short-term decision-makers.

### v3

@@ Abstract
The NBA draft can incentivize teams to deliberately lose. We propose a draft mechanism that is practical, incentive-compatible, and favors weaker teams. The Carry-Over Lottery Allocation (COLA) framework represents a paradigm shift in evaluating team quality, replacing single season standings with multi-year playoff outcomes. In our proposed mechanism, every non-playoff team receives the same number of lottery tickets, removing incentives to lose. Lottery tickets carry over to future lotteries, but playoff success or winning a top pick diminishes a team’s accumulated tickets. The lottery is familiar and preserves fan engagement.
Implementation challenges are addressed to demonstrate feasibility, including transitioning to COLA, handling trades, and accommodating draft classes of varying strength. For exceptionally strong classes, teams may prefer the lottery to the playoffs. We provide a solution, employing a truth-elicitation mechanism to identify such years and expanding lottery eligibility to include as many playoff teams as necessary to preserve incentive compatibility.
@@ Introduction
In the current NBA draft system, all non-playoff teams enter a lottery for the top four draft positions, with worse records giving higher chances of winning. Teams that do not win a top pick are then slotted by record, starting with the worst team. This structure creates a trade-off: the league wants to reward weaker teams to maintain competitive balance, but higher lottery odds for the worst teams also increase incentives to lose. Despite repeated adjustments to lottery probabilities over the decades, tanking persists because of a fundamental impossibility: Munro and Banchio [7] proved that no draft mechanism based solely on end-of-season records can simultaneously reward poor performance and eliminate incentives to tank. The current system is, mathematically, doomed to fail.
@@
We answer in the affirmative with the proposal of the Carry-Over Lottery Allocation (COLA) framework. COLA achieves incentive compatibility by eliminating any advantage from additional losses. Each team maintains a stockpile of tickets that grows and shrinks, but the stockpile never grows more because of additional losses. COLA rewards the weakest teams through carry-over: lottery tickets that do not win a top draft pick are retained for future lotteries, while playoff success or winning a top pick diminishes a team’s accumulated tickets.
COLA is a framework, and there are many possible variants within the framework. This paper discusses possible variants later, but primarily focuses on Classic COLA, which is one instance within the framework.
Classic COLA is practical to implement. It uses a familiar draft with lottery, requiring minimal changes to existing league procedures and fan expectations. Traded draft picks need special handling, but COLA solves this by excluding traded picks from lottery eligibility. The mechanism also accommodates variations in draft class strength. For a weak class, a team might view a lottery win as undesirable due to a loss of tickets in future lotteries, so there is an option to skip the lottery for a cost. For a strong draft class, many teams might prefer to be in the lottery instead of the playoffs. This violates a key assumption of anti-tanking mechanisms: reaching the playoffs is the primary goal, with draft picks as a secondary consideration. This problem has no solution in existing mechanisms, because if all teams prefer the lottery, no mechanism can both reward weak teams and prevent tanking. To address this, COLA identifies strong draft classes in advance using a Bayesian Truth Serum–based media survey and correspondingly expands lottery access as far as necessary but only when necessary, maintaining incentive compatibility without altering the core structure of the mechanism.
The remainder of this paper proceeds as follows. In Section 2, we review related work, including other draft mechanisms. Section 3 presents the core mechanism: how lottery tickets accumulate, how they are diminished by playoff success or lottery wins, and why this structure ensures that teams never maximize their expected utility by losing. Section 4 examines critical implementation challenges, including the handling of traded draft picks, transitioning from the current system, and addressing potential violations of our preference assumptions. Section 5 presents simulation results that provide further evidence that under COLA all teams have an equal likelihood of success over time. In an appendix, we provide a formal proof that there exists a draft mechanism that is incentive-compatible and favors the weakest team.
@@
  • Munro and Banchio [7] propose that the lottery be based on standings at a dynamically-determined point in the season. It meets the Preferencing and Anti-Tanking criteria under standard assumptions, but does not meet the Practicality criterion for two reasons. First, the calculation of the dynamically-determined date is opaque to a casual fan and requires knowledge of private team preferences. Choosing a predetermined early cutoff date may seem to approximate their approach, but if the date is announced and does not take into account teams’ incentives, then it simply moves tanking to earlier in the year. Second, in most seasons there are teams that believe their likelihood of playoff success is so small that they have an incentive to tank as soon as the season starts, in which case there are no standings on which to base the lottery.
  • The WNBA bases the lottery on the combined record of the two most recent seasons. This improves Preferencing for persistently weak teams but it is not an Anti-Tanking mechanism. A loss in one season could affect the draft position for two seasons, giving even playoff teams incentives to underperform.
  • Gold Plan [3] bases the draft on the number of wins that a team earns after playoff elimination. Under the assumption that every team puts forth maximum effort until they are mathematically eliminated, it satisfies Anti-Tanking. However, in practice it is common for teams to tank before they are eliminated. This mechanism can incentivize tanking earlier in the season so that they are eliminated as soon as possible. That way, the team has more time to rack up post-elimination wins, so due to practical concerns The Gold Plan is not truly Anti-Tanking. In terms of Preferencing, some weaker teams are favored over some stronger teams, but the Gold Plan disadvantages truly bad teams that cannot win many games even if they are eliminated early.
@@ A New Paradigm for Sports Drafts
With this paradigm shift, we can design a draft mechanism that meets all five of the desired criteria. The following section presents Classic COLA, a practical mechanism that implements these principles.
@@
We introduce the Carry-Over Lottery Allocation (COLA) framework, named for two of its key features: draft picks are allocated via lottery, and some lottery tickets carry over between years. While there are many possible variants in this framework, we will focus on one instance that we will call Classic COLA. For brevity, we refer to this simply as COLA throughout the paper, reserving the “Classic” distinction for contexts where it is necessary to differentiate it from other variants. We note that COLA thus refers both to the overall framework and to this specific instance, with the meaning clear from context.
@@
In the appendix, we provide a formal proof that, under reasonable assumptions, there exists a variant of COLA that is incentive-compatible and gives the weakest team the best odds of winning the first draft pick. Classic COLA satisfies the first property but not the second. Because it favors both weak and unlucky teams, the weakest team is not guaranteed to have the best odds of winning the first pick.
@@
Increment Size. Classic COLA uses an annual increment of $\alpha=1000$, but a different value could be used. To maintain incentive compatibility, it is important that a team not gain any benefit from additional losses, but it is not necessary for $\alpha$ to be constant. For example, the increment could be larger for teams that have missed the playoffs for multiple years in a row. It could also be based on other factors, even giving teams additional lottery tickets based on how many games they win. As long as losses do not earn more lottery tickets, incentive compatibility is maintained.
@@
Diminishment Rates. Classic COLA’s design (100%, 75%, 50%, 25% for draft picks; similar for playoff rounds) prioritizes simplicity. More aggressive rates emphasize recency; gentler rates give historical performance more weight. More aggressive rates on playoff success also increase the likelihood that the line must be moved to maintain incentive compatibility. More complex diminishment rules could be employed. Beyond percentages, diminishment could use fixed amounts, caps/floors, or escalation rules (e.g., doubled rates for teams making three consecutive playoff appearances).
@@
Lottery Scope. Following current NBA practice, the lottery for Classic COLA is for the top four picks, but this extends naturally to more or fewer positions if needed. The mechanism works identically with any number of lottery positions, though more positions increase the potential for the worst team to receive a poor pick despite having the best odds.
@@
Non-Lottery Pick Assignment. Picks 5-14 could be assigned by current-season standings (matching the existing NBA system) or by lottery index (consistent with COLA’s multi-year evaluation). Standings-based assignment provides familiarity, while lottery-index assignment rewards historically weak teams throughout the draft order, not just through the lottery. The latter also mitigates the incentive to tank for the purpose of retaining a traded draft pick that has protections. We recommend banning those sorts of trade protections if current-season standings are used to assign the non-lottery picks.
@@
No Lottery At All. If the draft pick order is determined by ordering all teams by highest lottery index (as suggested in the previous variant), then Playoff Track Record and prior drafts would determine the draft order instead of regular season standings. The increment rule, carry-over, and diminishment for playoff success and top draft picks accomplish their goals with or without a lottery. (Diminishment for top picks should remain, even if not determined by lottery. Otherwise, the same team would pick first each year until making the playoffs.) However, the lottery also generates fan interest, so preventing tanking is not its only purpose.
@@
Opt-Out Penalty. Classic COLA’s $2\alpha$ cost for skipping weak draft classes balances individual flexibility against the desire for league-wide participation. Lower the value to increase team autonomy. Adjust higher or eliminate the option if more participation is desired.
@@
Moving the Line. Our six-point scale on the analyst survey could be expanded to more options for more granularity at the expense of complexity for survey participants. For example, an alternative design could expand the survey to twelve options by allowing each playoff round to be included in the lottery either with or without diminishment for playoff success at that level.
Pros and cons for having the league office or a Blue Ribbon Commission decide how to set the line between lottery and non-lottery teams were discussed in Section 4.4. Alternatively, the line could be set as immovable. A static line is simpler, but it risks introducing a tanking hazard (if placed too low) or not providing enough help to weak teams (if placed too high).
@@
Increments in Strong Draft Years. When the line moves to include playoff teams, Classic COLA uses a binary rule: if eight or more teams remain excluded, all lottery participants receive $\alpha$; if four or fewer remain excluded, nobody receives $\alpha$. A simpler alternative would always grant $\alpha$ to all lottery participants regardless of where the line falls, prioritizing ease of explanation over maximizing Preferencing for Quality in strong draft years.
@@
Traded Picks Winning the Lottery. Under Classic COLA, traded draft picks are excluded from lottery eligibility. Alternative designs could allow traded picks to enter the lottery, with an appropriate lottery index diminishment rule. For example, diminishment could be applied to both the pick holder and the original team if the pick wins. This would introduce a third class of traded draft picks: protected 1–4, unprotected but lottery-ineligible, and unprotected lottery-eligible picks. While this variant improves trade-market liquidity and gives teams more flexibility, it is substantially more complex and creates the opportunity for systematic mismanagement of long-run lottery risk by short-term decision-makers.
@@ Conclusion
Appendix A Appendix: Proof of the Existence of a Draft Mechanism that is Incentive-Compatible and Favors the Worst Team
Munro and Banchio’s impossibility theorem leads us to ask whether, with a different definition of worst team, there exists an incentive-compatible draft mechanism that favors the worst team. Here, we prove that, under a reasonable definition of worst team and set of assumptions about team preferences, such a mechanism does exist.
League Structure
We consider a league with 30 teams and a standard playoff structure of 16 playoff teams and 4 playoff rounds. Seeding, conference differences, and home-court advantage are ignored. After the playoffs, there is a draft of 22 incoming players and a lottery to determine which teams will have the right to make those selections. (There may also be other draft picks beyond those 22, but they are assumed to have negligible value and are not factored into the model.)
Team Preferences and Definition of Worst Team
We define the worst team as the team with the longest active streak of seasons without winning a playoff series.
Each team maximizes an objective function with the following priorities:
  1. 1.
    Primary goal: playoff series wins $s$
  2. 2.
    Secondary goal: draft picks $d_{1}>d_{2}>\dots>d_{22}$
  3. 3.
    Tertiary goal: regular season game wins $g$
In other words, a team’s preferences satisfy:
$\text{Primary}\;>\;\text{Secondary}\;>\;\text{Tertiary}\;\gg\;\text{other considerations}.$
Let:
  • $x_{s}\in\{0,1,2,3,4\}$ denote the number of playoff series won,
  • $x_{k}\in\{0,1\}$ denote an indicator for receiving draft pick $k$,
  • $x_{g}\in\{0,1,\dots,82\}$ denote the number of regular-season games won.
Utility is given by
$U=x_{s}s+\sum_{k=1}^{22}x_{k}d_{k}+x_{g}g,$ (5)
where $s>d_{1}>d_{2}>\dots>d_{22}>g>0$.
For each game, a team chooses to either try to win or to lose. Games are independent, which means that if a team tries to win, tries to lose, does win, or does lose a game, it has no impact on the team’s ability to win a future game.
Diet COLA Mechanism
We define Diet COLA as a variant of COLA where each team $t$ has a lottery index $L_{t}$ that carries over from year to year.
  • If a team does not win a playoff series, then the team’s lottery index is unchanged for this year’s lottery. It will be incremented by one for the next year:
    $L_{t}^{\prime}=L_{t}+1.$ (6)
  • If a team wins a playoff series, then the team’s lottery index is reduced to zero:
    $L_{t}^{\prime}=0.$ (7)
All teams that fail to win a playoff series participate in the lottery, and 22 draft picks are distributed via lottery.
Each team $t$ that participates in the lottery has a number of tickets equal to $L_{t}$.
Draft pick trades are banned.
Pool Manipulation Assumption
The probability that team $t$ wins the first draft pick is:
$\text{P(\#1 pick)}=\frac{L_{t}}{\sum_{j\in\text{lottery}}L_{j}}.$ (8)
A team cannot affect $L_{t}$ other than the standard increment that occurs if they fail to win a playoff series. Since $s>d_{t}$, it is never better to lose in an attempt to increase $L_{t}$.
However, a team could attempt to change the denominator by manipulating the pool of lottery opponents. Suppose a team would lose to a high-index opponent, pushing that opponent into the playoffs, and then that high-index opponent would beat a low-index team that falls into the lottery. That would change the denominator, increasing the team’s probability of getting the number one pick.
There are several reasons why we do not expect teams to employ this technique.
  • Until it is late in the season, it is unknown whether or not a high-index team will end up near the playoff border where this sort of technique would even make a difference.
  • If a high-index team is successfully pushed into the playoffs, there is no guarantee that they win a playoff series, so they may end up in the lottery anyway.
  • If a high-index team does win a playoff series, the benefit is limited by the index of the team that replaces them.
  • If the technique is successful, the benefit of having a higher probability of getting a top pick is spread out among all the teams in the lottery, weighted toward the teams with the most tickets. Because of this, even a large swing in the size of the ticket pool means only a small gain for any individual team.
  • The nature of the draft mechanism will tend to minimize the maximum drought length, which also limits the potential benefit of this technique.
  • Normal tanking requires a team to be weak in general, but this technique requires losing a specific game, which is a different sort of intentional losing. This sort of targeted losing is akin to throwing a game for the purposes of gambling, and penalties for the latter could include a permanent ban from the league for anyone involved.
Simulations in Section 5 indicate that in 90% of seasons, this technique would not offer the opportunity for a significant probability increase. Under Diet COLA, where a team could lose in the playoffs but still be in the lottery, the technique would be even less effective.
Because of the limited and uncertain effectiveness of this technique and the potentially severe penalties if discovered, we make the reasonable assumption that teams value winning the game for its own sake more than the expected benefit of manipulating the pool of opponents in the lottery:
$g>\mathbb{E}[\text{gain from opponent manipulation}].$ (9)
Probability of Playoffs Wins and Lottery Participation
Note that if a team participates in the lottery, they will receive a draft pick that can be treated as a fixed value, $d_{t}$. Though the specific draft position is unknown until the lottery, teams are unable to affect which draft pick they will receive this year because they cannot improve $L_{t}$. Thus each team $t$ will either get draft pick with value $d_{t}$ or win a playoff series.
Let:
  • $P(s)$ = probability of eventually winning a playoff series
  • $P(d_{t})$ = probability of remaining in the lottery and receiving draft value $d_{t}$
Because these are complementary outcomes,
$P(s)+P(d_{t})=1.$ (10)
Thus,
$\Delta P(d_{t})=-\Delta P(s).$ (11)
Incentive Compatibility Proof
We compare the utility impact of winning versus losing a game.
Case 1: Before Elimination
If the team wins, the marginal utility is
$U_{W}$ $=s\,\Delta_{W}P(s)+d_{t}\,\Delta_{W}P(d_{t})+g$ (12)
$=s\,\Delta_{W}P(s)+d_{t}\left(-\,\Delta_{W}P(s)\right)+g$ (13)
$=\Delta_{W}P(s)\,(s-d_{t})+g,$ (14)
If the team loses:
$U_{L}$ $=s\,\Delta_{L}P(s)+d_{t}\,\Delta_{L}P(d_{t})+0$ (15)
$=s\,\Delta_{L}P(s)+d_{t}\left(-\,\Delta_{L}P(s)\right)$ (16)
$=\Delta_{L}P(s)\,(s-d_{t}).$ (17)
Because
$s>d_{t},\quad g>0,\quad\Delta_{W}P(s)>0,\quad\text{and}\quad\Delta_{L}P(s)<0,$
it follows that
$U_{W}>0\quad\text{and}\quad U_{L}<0.$
Therefore, it is strictly better to try to win.
Case 2: After Elimination
After elimination:
$P(s)=0,\quad P(d_{t})=1.$ (18)
Winning yields:
$U_{W}=g.$ (19)
Losing yields:
$U_{L}=0.$ (20)
Since $g>0$, winning is strictly preferred.
Case 3: After Clinching Playoffs
In this case:
$\Delta P(s)=0,\quad\Delta P(d_{t})=0.$ (21)
Winning yields $g$, losing yields 0, so again winning is strictly preferred.
Case 4: During the Playoffs
In the first playoff round:
$U_{W}$ $=\Delta_{W}P(s)(s-d_{t}),$ (22)
$U_{L}$ $=\Delta_{L}P(s)(s-d_{t}).$ (23)
Since $\Delta_{W}P(s)>0$ and $\Delta_{L}P(s)<0$, winning strictly dominates losing. (In this model, $s$ represents the inherent value of winning a playoff series, but no value has been assigned to individual playoff games. If a value were assigned to the individual games, it would not change the analysis.)
Main Result
Under Diet COLA, no team ever prefers to lose a game. The team with the longest active streak without a playoff series win has the highest lottery index. Thus, the highest lottery index implies the highest probability of receiving the first draft pick.
The model includes these assumptions:
  1. 1.
    Teams prefer to win a playoff series over getting any draft pick.
  2. 2.
    Teams prefer higher draft picks to lower draft picks.
  3. 3.
    Teams prefer any draft pick to winning any single game.
  4. 4.
    Teams prefer winning a game to other factors (which are considered in this model).
  5. 5.
    Teams prefer winning a game to the uncertain and limited benefit of attempting to manipulate the pool of opponents in the lottery.
Our main result is that under these assumptions, a simplified version of the Carry-Over Lottery Allocation draft mechanism (Diet COLA) is provably incentive-compatible, and under Diet COLA the team with the longest active streak of seasons without a playoff series win has the highest probability of receiving the first draft pick.
In other words, under reasonable assumptions, the Diet COLA draft mechanism is an anti-tanking draft mechanism that favors the worst team.